December 4, 2014 FDC+CDC meeting
- Instructions for Bluejeans meeting connection
- FDC meeting ID: 290664653
- To join via a Web Browser, go to the page  https://bluejeans.com/290664653.
- CDC analysis update (Beni, Mike)
- FDC analysis update (Lubomir, Vlad)
- Other: electronics, engineering
Participants: Sean (NU), Curtis, Naomi (CMU) Beni, Mike, Cody, Dave, Nick, Chris, Simon, Fernando, Vlad, and Lubomir (JLab).
- Most of the time we discussed the sagging of the pedestals at high rates as illustrated by Mike's picture linked above: for the straws from the inner rings (numbering of the straws on x starts from the inner tubes) we see lower pedestal values. Proposed explanation (Lubomir): negative tails especially at high (saturated) amplitudes can't recover before the next hit. Fernando: it is possible, according to  there are negative tails at the output of the pre-amp (examples of saturated signals not shown there) but it requires further studies together with the effects of the cable and fADC125 shaping. Fernando is going to study this on the bench.
- Beni showed examples of fADC pulses where you often see saturated signals followed by long positive tail, but the window is not wide enough to see what's next. We will try using the max number of samples for the trigger window, it is not clear what and where's the limit for that.
- Beni also demonstrated a code that calculates the pedestals every second. The above linked root file contains this information for runs 1512 to 1521.
- We discussed possible ways to study this issue that can be done during the high-luminosity CDC studies with the beam: lowering the HV, connecting a scope to one card, increasing the number of samples, comparing channels with high/low rate.
- Lubomir presented several slides for the run coordination meeting on the FDC performance (linked above): chamber efficiency ~95%, cathode resolution 150-200 microns for the third package that has been aligned in the EEL126.
- Vlad finished the first iteration of calculating the gain coefficients of all the cathodes (presentation linked): correlation b/n bad channels and gain coefficients, width of the gain coefficient distribution ~10%.
- Dave: the output pressure from the CO2 regulator was reduced to ~45 psi. After that the strange spikes seen in the pressure of the FDC tank (shown on the last meeting) disappeared.